OpenStudy (anonymous):

a=b 1=2 where is the fallacy?

OpenStudy (anonymous):

\[a^{2}=b^{2}\]

terenzreignz (terenzreignz):

<shrugs> What are we to do with this given equation?

OpenStudy (anonymous):

Prove it?

OpenStudy (anonymous):

let me change the equation

OpenStudy (unklerhaukus):

so many solutions

OpenStudy (unklerhaukus):

second line is false

terenzreignz (terenzreignz):

LOL... I think you're holding out on us XD

OpenStudy (anonymous):

hmm I'll make it longer

OpenStudy (unklerhaukus):

get a real question,

OpenStudy (anonymous):

a=b. a^2=ab a^2+a^2=a^2+ab 2a^2 wait I'm being called by my mom I'll continue this later http://www.math.toronto.edu/mathnet/falseProofs/first1eq2.html

OpenStudy (anonymous):

more simple: 1º: a=b 2º: ab=b^2 (multiply bouth sides by b) 3º: ab-a^2=b^2-a^2 (rest a^2 from bouth sides) 4º: a(b-a)=(b-a)(b+a) 5º: a=b+a 6º: a= a+a (since a=b) 7º: a=2a or 1=2 the falacy is in step 4º to 5º, because (b-a)=0. Remmember a=b. So we actualy devided by 0, which is not alowed

OpenStudy (anonymous):

it is, if a\(\neq =0\)

OpenStudy (anonymous):

what is your concern?

OpenStudy (anonymous):

i am curious, :)

OpenStudy (anonymous):

@Loser66 ?

OpenStudy (anonymous):

a=2a a/a=2a/a 1=2

OpenStudy (anonymous):

the problem is already with a=2a. But we got here because of step 4,5

OpenStudy (anonymous):

it is imposible to happen that a=2a for \(a \neq 0\)